Manager Chris Channing of Fabric Mills, Inc., has developed the forecast shown in the table for bolts of cloth. The figures are in hundreds of bolts. The department has a normal capacity of 275(00) bolts per month, except for the seventh month, when capacity will be 250(00) bolts. Normal output has a cost of $40 per hundred bolts. Workers can be assigned to other jobs if production is less than normal. The beginning inventory is zero bolts.

Month 1 2 3 4 5 6 7 Total
Forecast 250 300 250 300 280 275 270 1,925


a. Develop a chase plan that matches the forecast and compute the total cost of your plan. Overtime is $60 per hundred bolts. (Negative amounts should be indicated by a minus sign. Leave no cells blank - be certain to enter "0" wherever required. Omit the "$" sign in your response.)

Period 1 2 3 4 5 6 7 Total
Forecast 250 300 250 300 280 275 270 1,925
Output
Regular
Overtime
Subcontract
Output - Forecast
Inventory
Beginning
Ending
Average
Backlog
Costs:
Output
Regular $ $
Overtime
Subcontract
Inventory
Backorder
Total $ $


b. Would the total cost be less with regular production with no overtime, but using a subcontractor to handle the excess above normal capacity at a cost of $50 per hundred bolts? Backlogs are not allowed. The inventory carrying cost is $2 per hundred bolts. (Round your Average valuesto 1 decimal place. Negative amounts should be indicated by a minus sign. Leave no cells blank - be certain to enter "0" wherever required. Omit the "$" sign in your response.)

Period 1 2 3 4 5 6 7 Total
Forecast 250 300 250 300 280 275 270 1,925
Output
Regular
Overtime
Subcontract
Output - Forecast
Inventory
Beginning
Ending
Average
Backlog
Costs:
Regular $ $
Overtime
Subcontract
Inventory
Backorder
Total $ $
Answer #1

answer:

a) Under pursue plan, production should be coordinated to demand forecast every month, subject to the condition that normal capacity limited is restricted to 275(00) bolts every month and 250(00) in the last month. So, in the table below, period 1 creation will be equivalent to period 1 demand,i.e, 250(00) bolts (Which is within normal capacity of 275(00) ). In period 2, demand is 300 but normal capacity is is restricted to 275, so regular production is 275 and balance 300 - 275 = 25 to be produced utilizing additional time. Additionally, for additional periods. In period 7, request is 270, in any case, normal capacity is 250, so again 20 will be created utilizing extra time.

Further, inventory will be 0 at all times since starting inventory is 0 and for every period creation is coordinated to demand forecast.

Likewise, output costs are determined as regular cost being regular output during the period * $40 per hundred bolts, so for instance, for period 1, regular output cost = 250 * $40 = $10,000 and correspondingly, for periods 2 to 7. Total regular production cost = $74,000.

Extra time costs for every period are determined as additional time production in every period * $60. Thus, for period 2, additional time cost = 25 * $60 = $1,500. Total overtime cost = $4,500

Subcontract, inventory and backorder costs are nil.

Period 1 2 3 4 5 6 7 Total
Forecast 250 300 250 300 280 275 270 1925
Output
Regular 250 275 250 275 275 275 250 1850
Overtime 0 25 0 25 5 0 20 75
Subcontract 0 0 0 0 0 0 0 0
Output-forecast
Inventory
Beginning 0 0 0 0 0 0 0 0
Ending 0 0 0 0 0 0 0 0
Average 0 0 0 0 0 0 0 0
Backlog 0 0 0 0 0 0 0 0
Costs
Output
Regular 10000 11000 10000 11000 11000 11000 10000 74000
Overtime 0 1500 0 1500 300 0 1200 4500
Subcontract 0 0 0 0 0 0 0 0
Inventory 0 0 0 0 0 0 0 0
Backorder 0 0 0 0 0 0 0 0
Total 10000 12500 10000 12500 11300 11000 11200 78500

b)  Under regular production methodology, normal capacity output is delivered in every month. If same is less than demand, subcontracting alternative is utilized and if same is in excess of demand, inventory is put away. In this way, in the table below, for period 1, normal capacity output is 275 while demand is 250, so abundance (inventory) is 25. Next, in period 2, starting inventory= period 1 consummation inventory = 25, production = 275, and demand= 300, so finishing inventory= 25+275-300 = 0. Normal inventory for every period is determined as (finishing inventory of current period + finishing inventories of prior periods)/No. of periods including current period. In this way, for period 2, normal inventory is (25+0)/2 = 12.5. In period 3, normal inventory is (25+0+25)/3, etc. We see the requirement for subcontract arises in period 5, when production less than demand and there is no stock to meet the abundance. In this way, subcontracting for 300 - 275 = 5(00) bolts.

Next, normal production cost is $40 * every period's regular production.

Subcontracting cost is $50 * every period's subcontracted production

Inventory conveying cost is $2 * finishing inventory (inventory conveyed) in every period.

Including everything gives us total cost as $77,350, which is lower than the total cost in choice.

Period 1 2 3 4 5 6 7 Total
Forecast 250 300 250 300 280 275 270 1925
Output
Regular 275 275 275 275 275 275 275 1900
Overtime 0 0 0 0 0 0 0 0
Subcontract 0 0 0 0 5 0 20 25
Output-forecast
Inventory
Beginning 0 25 0 25 0 0 0 0
Ending 25 0 25 0 0 0 0 0
Average 25.0 12.5 16.7 12.5 10.0 8.3 7.1 92.1
Backlog 0 0 0 0 0 0 0 0
Costs:
Output
Regular 11000 11000 11000 11000 11000 11000 10000 76000
Overtime 0 0 0 0 0 0 0 0
Subcontract 0 0 0 0 250 0 1000 1250
Inventory 50 0 50 0 0 0 0 100
Backorder 0 0 0 0 0 0 0 0
Total 11050 11000 11050 11000 11250 11000 11000 77350